OpenStudy (anonymous):

find the equation of the line. Write the equation using function notation. Through (-4,-3); perpendicular to f(x)=-3x+3

OpenStudy (campbell_st):

gradient of f(x) is m = -3 the perpendicular gradient is -1/m or -1/-3 then m = 1/3 using point gradient formula y +3 = 1/3(x +4) y = 1/3x - 5/3 then f(x) = 1/3x - 5/3